UARS Weekly Status Report

06 December 2002


                UARS WEEKLY STATUS REPORT - 06 Dec 2002



Other than being off for the solar eclipse this week, SOLSTICE continues
to operate normally, taking solar data on all available orbital tracking


After resuming operations on Monday, Dec. 2, SUSIM was again
switched off for the eclipse.  Otherwise, SUSIM successfully
gathered all of its daily mid-resolution and MgII solar scans
this past week during periods of predicted UARS telemetry.


Nothing received.


     The PEM web site has moved.  It has a new URL.  Access the PEM web site
using  The PEM data from the GUVI conjunction
periods has moved to  It has been
discovered that this web site was not updated properly with the latest PEM
data plots.  SwRI will investigate the problem.

     Level 2 PEM HEPSA data files are being reprocessed for submission to the 
DAAC.  This data has been designated as version 2; however, there is no format
difference from the version 1 data format.  Version 2 HEPSA data is being 
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files.  Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files.  Currently, HEPSA data between 1991274 and 1993365 has been 
reprocessed and is archived in the DAAC.  HEPSA data reprocessing continues
but has been slowed temporarily while the first 30 days of 1994 data is being 

     PEM continues to work with the GUVI team on the diagnosis of data from
the first conjunction period.  Analysis of the two conjunction passes showed
that the aurora as detected by GUVI was at higher latitude than UARS.  
GUVI images from other times from the first conjunction period show spatial
regions that both GUVI and PEM sampled when aurora was present.  Separation
is several minutes to 10's of minutes when both instruments saw aurora. 
PEM has seen no aurora during the second conjunction period.  Due to a 
spacecraft anomaly, no PEM data was collected during the third conjunction 

     A poster presentation on the GUVI-PEM comparison from the first conjunction
period is being presented at the American Geophysical Union Fall Meeting.
The poster is titled "Validation of the GUVI Auroral Radiance Measurements 
Using the UARS/PEM Instrument" by  A Frahm, G Crowley, L J Paxton, C H Jackman,
A B Christensen, D J Strickland, D Morrison, Y Zhang, H Kil, B C Wolven, 
J D Winningham, J R Sharber, S K Avery, J D Craven, R R Meier, C I Meng, 
P R Straus, C Swenson, and R Waltersheid.  An additional presentation on the
overall storm period titled "The Relative Atmospheric Impacts and Energy 
Inputs of Precipitating Solar and Magnetospheric Ion and Electron Populations 
during the 17-24 April 2002 Events" is being presented by J U Kozyra, 
D N Baker, G Crowley, D S Evans, X Fang, R A Frahm, S G Kanekal, M W Liemohn, 
G Lu, G M Mason, R A Mewaldt, L J Paxton, E C Roelof, and J D Winningham at the
same meeting.

     PEM continues to receive NMC data files from NOAA.  No problems have
been encountered.  NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.

*Beginning of power orbit day (PEM activation): array current >5 amps, array 
     voltage >53.5 volts, and battery current >2.0 amps.  
 End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
     array current <9 amps, and battery currents <0.2 amps.
 Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.


Nothing received.


Nothing received.


Nothing received.


Nothing received.


The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.

Daily processing of instrument Daily Activity Plans continues
with no problems.


Nothing received.
                           REPORTING PERIOD

This report covers 23 November 2002 (Orbit 61256, GMT Day 327) through
29 November 2002 (Orbit 61360, GMT Day 333).

                         SPACECRAFT OPERATION

The observatory is now in Forward Flight (Southern Hemisphere viewing)
and   is  performing  nominally  under  two-battery  operations.   The
instrument operational changes were:


    Instrument    Time          Orbit     Comment
    HALOE         328/1203      61278     POWER ON
    ACRIM         329/1715      61296     POWER ON QB A
    ACRIM         330/0104      61301     POWER ON QB B ON / QB A OFF
    WINDII        330/1303      61303     POWER ON
    ACRIM         330/1452      61310     POWER ON QB A ON / QB B OFF
    ACRIM         331/0138      61316     POWER ON QB B ON / QB A OFF
    ACRIM         331/1303      61324     POWER ON QB A ON / QB B OFF
    ACRIM         332/0136      61331     POWER ON QB B ON / QB A OFF
    ACRIM         332/1302      61339     POWER ON QB A ON / QB B OFF
    PEM           332/1339      61339     AXIS-1 POWER ON
    ACRIM         333/0134      61346     POWER ON QB B ON / QB A OFF
    ACRIM         333/1301      61354     POWER ON QB A ON / QB B OFF

The  beta  angle  ranged between the angles of 14.3  degrees  to  31.4
degrees  for  this  report period.  The beta angle is  now  increasing
towards a maximum angle of 35 degrees on 04 December 2002 (DOY 338).

Spacecraft  battery  1,  2,  and 3 performance  monitoring  continues.
Battery 1 remains OFF the charge relay.  Battery 2 and Battery  3  are
currently maintaining greater than 23.6 V end-of-night (EON) load  bus
voltage.   Battery  temperatures are stable with a  temperature  delta
between Battery 2 and Battery 3 of 1.27 to 1.29 degrees C.  Additional
battery  performance  data for each battery  on  22  November  and  29
November 2002 are provided in Appendix D.

The  clock  error  ranged  from -8.2 to 0.8 msec  during  this  report
period.  Clock rate adjustments are listed in Appendix C.

All  instrument  and subsystem engineers were advised on  the  initial
excursion  of ALERTS or Out-Of-Limits conditions detected by  the  FOT
for  this report.  A summary is listed below and the details of  ALERT
and Out-Of-Limits occurrences are listed in Appendix B.

    OUT-OF-LIMITS  =    00
    ALERTS         =    00

The  following real-time command sheets were executed for  the  listed
operational element(s) during this report period (see Appendix A).

    FOT       =    01
    HALOE     =    02
    WINDII    =    01
    PEM       =    01
    ACRIM     =    02

UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.

    FOT  =    00

One  (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).

                        GROUND SYSTEM OPERATION

Zero  (0) operational support problems occurred this period.   Anomaly
Reports are included in Appendix E.

The  data  loss  calculations are no longer  being  supplied  by  Data
Capture Facility (DCF).

                       OTHER SIGNIFICANT EVENTS

                       UARS SATELLITE OPERATIONS

A new TAM alignment matrix was uplinked on 21 November 2002 to improve
the attitude performance.  As of last week there was a noticeable
difference but with a highly converged filter the attitude error are
taken out very slowly, so the FOT wasn't sure whether the new
alignment had fixed the problem.  After monitoring the attitude over
the course of the next week the FOT observed much lower FSS residuals
(on the order of 100 arc seconds) and the attitude seems to be
behaving very well.  The FOT is now confident in the attitude solution
for forward flight and will be ready to closely monitor the reverse
flight behavior after the yaw maneuver on 18 December 2002.

The FOT is also in the process of preparing for the upcoming Solar
Eclipse on 04 December 2002.  The analysis provided by Flight Dynamics
shows that there will be a significant impact to UARS' power during
the eclipse.  The eclipse will take away available solar energy for 3
consecutive orbits.  The worst of the three orbits shows a 19 minute
period during UARS power day with up to 66% shadowing.  This impact to
UARS power forces the FOT to put all instruments in their off or
standby modes until at least one or two orbits following the eclipse.
The eclipse will begin effecting UARS at 338/0522 and last until
338/0928.  The instruments affected by the eclipse are HALOE,


The FOT attended a NASA Anomaly Review Meeting to discuss the OBC halt
anomaly experienced on 21 October 2002.

                       FUTURE SIGNIFICANT EVENTS

There  will be a Solar Eclipse on 04 December 2002 (DOY 338) beginning
at approximately 0500 until 1000 UTC.

The  next  UARS Yaw slew is scheduled for 18 December 2002 (DOY  352).
This  maneuver will be a Forward to Reverse Flight maneuver  (Northern
Hemisphere viewing).

Attachments:   Appendix A,     Real-time Command Sheets
               Appendix B,     Out-Of-Limits   &   ALERT Occurrences
               Appendix C,     Clock &  Solar  Array  Rate Corrections
               Appendix D,     UARS Battery Performance Data
               Appendix E,     UARS Anomaly Reports

                              APPENDIX A
                       REAL-TIME COMMAND SHEETS
                     Week ending 29 November 2002

ORBIT   AOS        NO.      SUBSYS      REASON
-----   ---        ---      ------      ------
61278   328/1203   02-459   HALOE       POWER ON (P-1)
61281   328/1717   02-460   HALOE       POWER ON (P-2)
61296   329/1715   02-461   ACRIM       POWER ON QB A
61301   330/0104   02-462   ACRIM       POWER ON QB B ON / QB A OFF
61309   330/1303   02-463   WINDII      POWER ON
61310   330/1452   02-461   ACRIM       POWER ON QB A ON / QB B OFF
61313   330/1955   02-464   CDH         PMP-B TO MODE F
61316   331/0138   02-462   ACRIM       POWER ON QB B ON / QB A OFF
61324   331/1303   02-461   ACRIM       POWER ON QB A ON / QB B OFF
61331   332/0136   02-462   ACRIM       POWER ON QB B ON / QB A OFF
61339   332/1302   02-461   ACRIM       POWER ON QB A ON / QB B OFF
61339   332/1339   02-465   PEM         AXIS-1 POWER ON
61346   333/0134   02-462   ACRIM       POWER ON QB B ON / QB A OFF
61354   333/1301   02-461   ACRIM       POWER ON QB A ON / QB B OFF

                              APPENDIX B
                     Week ending 29 November 2002

The following Out-Of-Limits occurrences were detected:

--------       -----   -----------                ------
-None this report period.

The following ALERT occurrences were detected:

MNEMONIC     DESCRIPTION                          ORBITS
--------     -----------                          ------
-None this report period.

                              APPENDIX C
                     Week ending 29 November 2002


ORBIT          TIME                   CLOCK RATE CHANGE
-----          -----                  -----------------
-None this report period.

                              APPENDIX D
                     Week ending 29 November 2002

     22 November 2002 - GMT Day 326 Beta = 10.8 deg, SA Pos = 269 deg

     DIFF V (mV)  TEMP (C)     CURR (amp)  EON V    AVE C/D  AVE
BAT  MAX/MIN      MAX/MIN      MAX/MIN     MIN      RATIO    DOD(%)
---  -------      --------     ----------  -----    -------  ------
1  -156.8/-179.2 -3.68/-4.94  +00.0/+00.0  08.3     0.000     0.0
2   +11.2/-22.4  +4.51/+3.23  +28.8/-13.2  26.2     1.040    19.5
3   +28.0/-22.4  +3.23/+2.29  +28.0/-12.8  26.2     1.020    19.1

     29 November 2002 - GMT Day 333 Beta = 31.4 deg, SA Pos = 269 deg

     DIFF V (mV)  TEMP (C)     CURR (amp)  EON V    AVE C/D  AVE
BAT  MAX/MIN      MAX/MIN      MAX/MIN     MIN      RATIO    DOD(%)
---  -------      --------     ----------  -----    -------  ------
1  -156.8/-179.2 -3.05/-3.99  +00.0/+00.0  08.4     0.000     0.0
2   +05.6/-123.2 +6.43/+5.14  +28.8/-15.2  25.0     1.040    22.0
3   +22.4/-67.2  +5.14/+3.87  +28.4/-16.0  24.8     1.030    21.5

ORB#           AOS TIME   LVL FR-TO
-----          --------   ---------
- None this report period.

                              APPENDIX E
                         UARS ANOMALY REPORTS
                     Week ending 29 November 2002

------    ------    -----     ----      ---       -----------------
-None this report period.

------    ------    -----     ----      ---       -----------------
02-057    FOT       60630     285/1008  30382     GENERIC LATE ACQUISITION
02-059    FOT       61191     322/1646  31767     YAW LATE ACQ (HGA TRACK)

------    ------    -----     ----      ---       -----------------
-None this report period.

                         OPEN ANOMALY REPORTS
------    ------    -----     ----      ---       -----------------
02-058    FOT       61113     317/1341  N/A       NEG ACQ / UFLTR ERROR